For a class field trip, the school would have to pay $750 for the bus for everyone signed up now. if 10 more students sign up, the price per person would go down by $2.50. How many students are signed up now?
Answer by richwmiller(17219)

x*y=750
(x+10)(y-2.50)=750
There were 50 students paying 15 each
There are now 60 students paying 12.50 each